OpenStudy (anonymous):

Consider the area shown below. The curve drawn is x^2 + y^2 = 3, and we have used the notation Dy for \Delta y.

OpenStudy (anonymous):

i know how to make a reiman sum, but after that for writing an integral, how do you go about finding the bounds

OpenStudy (aum):

This is quarter of a circle and the radius of the circle is \(\sqrt{3}\). Limits are x = 0 to x = \(\sqrt{3}\).

OpenStudy (anonymous):

\[so its \int\limits_{a}^{b} \sqrt{3-y^2}dy\]

OpenStudy (aum):

y = 0 to y = \(\sqrt{3}\)

OpenStudy (anonymous):

how wouuld you go about setting up an equation to find those? do you just substitute values in...?

OpenStudy (aum):

The uppermost point on the curve is on the y-axis where x = 0 put x = 0 and find y.

OpenStudy (anonymous):

oh then when it hits the x axis y=0 and that gives you sqrt3 for the b